If X & Y are 2 x 2 matrices with integer elements such that X3+Y3 = A, then find the det (X + Y)

Grade:12th pass

1 Answers

Sher Mohammad IIT Delhi
9 years ago
given X & Y are 2 x 2 matrices with integer elements such that X3+Y3= A, by intuition and checing conditions
x=[1 0; 0 0]
y=[0 0;0 1]
x+y=[1 0;0 1]
det(x+y)=1